# KVCompose: Efficient Structured KV Cache Compression with Composite Tokens

### v1

@@ 3.2 Attention-Based Token Importance Scoring
Our approach leverages attention patterns to identify which context tokens are most critical to retain during KV cache compression. The central hypothesis is that tokens receiving consistently high attention across tasks are more influential for preserving model behavior and therefore should be prioritized during retention. Consequently, we estimate token importance by analyzing aggregated attention scores on a representative task set $\mathcal{T}$, which guides cache eviction decisions.
@@ 3.2.1 Task Set Construction
The effectiveness of the proposed scoring procedure depends on the construction of the set of tasks $\mathcal{T}$ used to estimate the importance of the token. We consider three practical settings:
@@ 3.3 Layer-wise Compression Strategy with Composite Tokens
In practice, KV caches are stored as tensors where all heads within the same layer share a common sequence dimension. This tensorized representation enforces that the number of retained elements must be identical across heads in a given layer. However, a key observation follows from the self-attention formulation in equation 2–equation 4: although the number of retained elements must be uniform across heads, there is no requirement that these correspond to the same original context tokens across heads. This insight enables a more flexible compression scheme through composite token construction. Rather than selecting a common subset of context tokens for all heads in a layer, we allow each head to independently select its most important elements, which are then aligned into shared positions (see Figure 1). We refer to these new tokens as the composite tokens.
@@ 4.2 Experimental Setup
In our experiments, $r$ varies in $\{0,0.1,0.25,0.4,0.5,0.6,0.7,0.8,0.9\}$. Moreover, we use max-pooling for $\text{Agg}_{\text{task}}$ in equation 14 and average-pooling for all other aggregation operations.
Table 2: Performance comparison of structured KV cache compression strategies under a predefined tolerance $\epsilon_{0}$.
@@ 4.3 Key Results
Table 3 compares methods using the AUC metric, which summarizes robustness across all compression ratios. KVCompose achieves the highest AUC on five out of six evaluation settings and the best overall average ($82.3$). The improvements over strong structured baselines are substantial: +8.9 points over TOVA and +15.5 points over PyramidKV. These results confirm that KVCompose degrades more gracefully as compression increases, preserving accuracy even in high-compression regimes.
@@
Beyond accuracy gains, KVCompose respects the tensorized cache structure required by existing inference engines, unlike methods that depend on per-head variability (e.g., SnapKV, KVzip). This ensures that the observed improvements translate directly into real memory and compute savings without requiring custom kernels or engine modifications. While our primary focus is on structured eviction methods, we also examine an unstructured variant of our approach and compare it against unstructured KV cache compression strategies. The results are reported in Appendix A.
Table 3: AUC comparison of structured KV cache compression strategies.
@@ Appendix A Appendix
In this appendix, we evaluate an unstructured variant of our approach, obtained by removing the final aggregation step and allowing each head to select its tokens independently. For completeness, we compare the unstructured variant of our approach with unstructured KV cache compression strategies, summarized in Table 4. Unlike structured eviction, these approaches assign nonuniform budgets across heads and evict individual entries independently. Although this often yields higher accuracy due to increased flexibility, such methods are not directly compatible with standard KV cache implementations (Jegou et al., 2024) and therefore require nontrivial engine modifications.

### v2

@@ 3.2 Attention-Based Token Importance Scoring
Our approach leverages attention patterns to identify which context tokens are most critical to retain during KV cache compression. The central hypothesis is that tokens receiving consistently high attention across tasks are more influential for preserving model behavior, and therefore should be prioritized during retention. Accordingly, we estimate token importance by analyzing aggregated attention scores over a representative task set $\mathcal{T}$, which guides cache eviction decisions.
@@ 3.2.1 Task Set Construction
The effectiveness of the proposed scoring procedure depends on the construction of the task set $\mathcal{T}$ used to estimate token importance. We consider three practical settings:
@@ 3.3 Layer-wise Compression Strategy with Composite Tokens
In practice, KV caches are stored as tensors where all heads within the same layer share a common sequence dimension. This tensorized representation enforces that the number of retained elements must be identical across heads in a given layer. However, a key observation follows from the self-attention formulation in equation 2–equation 4: although the number of retained elements must be uniform across heads, there is no requirement that these correspond to the same original context tokens across heads.
This insight enables a more flexible compression scheme through composite token construction. Rather than selecting a common subset of context tokens for all heads in a layer, we allow each head to independently select its most important elements, which are then aligned into shared positions (see Figure 1). We refer to these new tokens as the composite tokens.
@@ 4.2 Experimental Setup
In our experiments, $r$ varies in $\{0,0.1,0.25,0.4,0.5,0.6,0.7,0.8,0.9\}$. Moreover, we use max-pooling for $\text{Agg}_{\text{task}}$ in equation 14 and average-pooling for all other aggregation operations. In Appendix A.1, we explore how different aggregation choices affect performance; we also perform ablation studies on some key aspects of the presented method.
Table 2: Performance comparison of the maximum compression ratio achieved by different structured KV cache compression strategies under a predefined tolerance $\epsilon_{0}$.
@@ 4.3 Key Results
Table 3 compares methods using the AUC metric, which summarizes robustness across all compression ratios. KVCompose achieves the highest AUC on five out of six evaluation settings and the best overall average ($82.3$). The improvements over strong structured baselines are substantial: +8.9 points over TOVA and +15.5 points over PyramidKV. These results confirm that KVCompose degrades more gracefully as compression increases, preserving accuracy even in high-compression regimes. More granular results can be found in Appendix A.3.
@@
Beyond accuracy gains, KVCompose respects the tensorized cache structure required by existing inference engines, unlike methods that depend on per-head variability (e.g., SnapKV, KVzip). This ensures that the observed improvements translate directly into real memory and compute savings without requiring custom kernels or engine modifications. While our primary focus is on structured eviction methods, we also examine an unstructured variant of our approach and compare it against unstructured KV cache compression strategies. The results are reported in Appendix A.2.
Table 3: Performance comparison of the AUC achieved by different structured KV cache compression strategies.
@@ Appendix A Appendix
In this section, we present additional experiments to further assess the performance of the proposed strategy for KV cache compression.
A.1 Ablation studies
To better understand the contribution of each design component, we conduct ablation studies on Llama3.1-8B and Qwen2.5-7B-Instruct using $10\%$ of data randomly sampled from the Ruler-4096 benchmark.
A.1.1 Effect of aggregation operators
Figure 2: Effect of aggregation operators on the model accuracy in task-agnostic and task-aware scenarios. We use the label $\text{Agg}(x,y,z),~x,y,z\in\{\text{avg},\text{max}\}$ to denote a configuration where $\text{Agg}_{\text{task}}=x$, $\text{Agg}_{\text{group}}=y$, and $\text{Agg}_{\text{head}}=z$. Results are reported for Llama-3.1 8B.
In Figure 2, we report results comparing different operator choices, i.e., max-pooling and average-pooling, for $\text{Agg}_{\text{task}}$, $\text{Agg}_{\text{group}}$, and $\text{Agg}_{\text{head}}$ To maintain a good level of readability, we report in Figure 2 the most significant operator combinations. Max-pooling for $\text{Agg}_{\text{task}}$ consistently yields stronger performance, indicating that preserving the strongest attention signal is most informative for token importance estimation. Keeping this setting, we also observed that for group and head aggregation, average-pooling provides the best accuracy for task-agnostic scenarios, while max-pooling leads to slight improvements in task-aware scenarios.
A.1.2 Effect of the mean on the token scores
In this appendix, we assess the benefit of augmenting token scores with their mean across all heads (see equation 16). Figure 3 shows that this augmentation (W/ mean) improves performance across compression ratios, in particular when using Qwen2.5-7B-Instruct in task-aware scenarios. Without the mean term (W/o mean), token importance estimates vary more strongly across heads, which can lead to inconsistent selections and accuracy loss under compression. Adding the mean provides a stabilizing signal, resulting in more reliable retention decisions and higher overall accuracy.
Figure 3: Effect of the mean on the token scores. Results are reported for Qwen2.5-7B-Instruct in task-aware setting.
A.1.3 Effect of Value Norms on Token Scores
Figure 4: Effect of incorporating value norms into token scores on the model accuracy. Results are reported for task-agnostic settings.
ExpectedAttention (Jegou et al., 2024) has shown that scoring tokens based on attention weights only has some drawbacks, as this approach ignores the scale of the value vectors. Indeed, two tokens with identical attention weights may contribute very differently if their value vectors have significantly different norms. Recalling equation 4, the output of an attention layer can be rewritten as:
$\mathbf{y}$ $=\sum_{i=1}^{H_{\text{q}}}\mathbf{A}_{i}\mathbf{V}_{i}\mathbf{W}_{i}^{O}$ (25)
$=\sum_{i=1}^{H_{\text{q}}}\big(\mathbf{A}_{i}\|\mathbf{V}_{i}\|\big)\,\frac{\mathbf{V}_{i}}{\|\mathbf{V}_{i}\|}\mathbf{W}_{i}^{O}$ (26)
$=\sum_{i=1}^{H_{\text{q}}}\big(\mathbf{A}_{i}\|\mathbf{V}_{i}\mathbf{W}_{i}^{O}\|\big)\,\frac{\mathbf{V}_{i}}{\|\mathbf{V}_{i}\mathbf{W}_{i}^{O}\|}\mathbf{W}_{i}^{O}.$ (27)
Equations 26 and 27 show explicitly that the effective contribution of a token is scaled by its value norm (either before or after projection). Empirically, we observe that token value norms can vary by nearly two orders of magnitude (approximately 11-$100\times$), indicating that norm differences are non-negligible in practice.
Motivated by this observation, we experiment with token scoring schemes that weight attention by the value norm. Specifically, we define a weighted attention matrix $\mathbf{A}_{i}^{(w)}$ in which each entry is multiplied by either the raw value norm $\|\mathbf{V}_{i}\|$ (v-norm) or the projected norm $\|\mathbf{V}_{i}\mathbf{W}_{i}^{O}\|$ (vo-norm). The v-norm variant corresponds to the approach used in ExpectedAttention (Jegou et al., 2024).
Our experiments reveal that the impact of the weighted attention matrix is model-dependent: we observed modest gains for Qwen models in some configurations, but a clear degradation for LLama-3.1 8B (see Figure 4). These results show that the attention-only criterion is a reliable and simple approach for our composite token scoring mechanism.
A.2 Discussion on unstructured eviction methods
@@
In this appendix, we evaluate an unstructured variant of our approach, obtained by removing the final aggregation step and allowing each head to select its tokens independently. For completeness, we compare the unstructured variant of our approach with unstructured KV cache compression strategies, summarized in Table 4. Unlike structured eviction, these approaches assign non-uniform budgets across heads and evict individual entries independently. While this often yields higher accuracy due to increased flexibility, such methods are not directly compatible with standard KV cache implementations and therefore require non-trivial engine modifications (Feng et al., 2024; Kim et al., 2025).
@@
A.3 Additional Performance Evaluation
In this appendix, we provide detailed results on the performance reported in Section 4.3. Specifically, Figure 5 presents a detailed comparison of the average performance of our method against state-of-the-art structured eviction strategies on the Ruler-4096 dataset (Hsieh et al., 2024). Moreover, Figure 6-Figure 11 further breaks down the results of Figure 5 at the task level, providing a fine-grained view of performance across the $13$ tasks in Ruler-4096.
Remark: PyramidKV extends SnapKV by assigning a predefined per-layer budget, giving larger budgets to lower layers. As the compression ratio increases, PyramidKV converges to SnapKV, which explains their identical performance in the high-compression regime ($r\geq 70\%$). In the mid-compression range ($r\approx 50\%$), however, the fixed budgeting scheme does not appear to align well with all models. This is particularly evident for Qwen2.5-7B-Instruct and Qwen3-14B-Instruct, where performance occasionally drops below that observed at higher compression ratios, producing an unusual curve shape.
Figure 5: Performance comparison of different structured eviction methods for different compression ratios.
Figure 6: Performance comparison of the different eviction methods per sub-tasks of the Ruler dataset. Results are reported for LLaMA-3.1-8B in the task-agnostic setting.
Figure 7: Performance comparison of the different eviction methods per sub-tasks of the Ruler dataset. Results are reported for Qwen2.5-7B-Instruct in the task-agnostic setting.
Figure 8: Performance comparison of the different eviction methods per sub-tasks of the Ruler dataset. Results are reported for Qwen3-14B-Instruct in the task-agnostic setting.
Figure 9: Performance comparison of the different eviction methods per sub-tasks of the Ruler dataset. Results are reported for LLaMA-3.1-8B in the task-aware setting.
Figure 10: Performance comparison of the different eviction methods per sub-tasks of the Ruler dataset. Results are reported for Qwen2.5-7B-Instruct in the task-aware setting.
Figure 11: Performance comparison of the different eviction methods per sub-tasks of the Ruler dataset. Results are reported for Qwen3-14B-Instruct in the task-aware setting.